jack is building a rectangular dog pen that he wishes to enclose. the width of the pen is 2 yards less than the length. if the area of the of the dog pen is 15 square yards, how many yards of fencing would he need to completely enclose the pen?
Answer by checkley77(12569)

W=L-2
L*(L-2)=15
L^2-2L-15=0
(L-5)(L+3)=0
L-5=0
L=5 ANSWER FOR THE LENGTH
W=5-2=3 ANSWER FOR THE WIDTH.
PROOF:
5*3=15
15=15
THEREFORE THE PERIMETER OF THIS PEN IS:
2*3+2*5
6+10=16 YARDS OF FENCING IS NEEDED.